OpenStudy (anonymous):

(5x+7y-12)^2 + (3x+2y-5)^2 =0. Find 9x+11y

OpenStudy (anonymous):

Hello bajji :) u have something like this\[a^2+b^2=0\]sum of 2 non-negative thing is zero so both of them must be zero\[(5x+7y-12)^2 + (3x+2y-5)^2 =0\]gives\[5x+7y-12=0\]\[3x+2y-5=0\]this is a simple system of equations, solve it by elimination or other methods which u are comfortable with

OpenStudy (anonymous):

Thanks a lot!!!! :D

OpenStudy (anonymous):

yw :)